OpenStudy (anonymous):

Need help with a complex numbers problem. Write -7+4i in trigonometric form. I know that R = sqrt 65 and theta = 0.519 ( Due to Arctan -4/7) but I don't know how to get the values of sine and cosine.

OpenStudy (anonymous):

you don't get the values of sine and cosine, you have them, they are \(-7\) for cosine and \(4\) for sine what you need is \(r\) and \(\theta\)

OpenStudy (anonymous):

unfortunately \(\theta\) is not \(\tan^{-1}(-\frac{4}{7})\) because you are in quadrant II

OpenStudy (anonymous):

\[-7+4i=r\left(\cos(\theta)+i\sin(\theta)\right)\]

OpenStudy (anonymous):

\[r\cos(\theta)=-7, r\sin(\theta)=4\] i made a mistake when i said \(\sin(\theta)=4\) etc, of course that is not possible. what i meant to say was \(r\sin(\theta)=4\)

OpenStudy (anonymous):

your \(r\) is correct, it is \(\sqrt{7^2+4^2}=\sqrt{65}\) but your \(\theta\) is wrong

OpenStudy (anonymous):

How would you get theta in this situation? I got Tan(theta) = b/a = 4/-7 then I used arc tan to get my answer which gives me -.519146

OpenStudy (anonymous):

are you working in degrees or radians?

OpenStudy (anonymous):

Radians

OpenStudy (anonymous):

you can only use \(\theta=\tan^{-1}(\frac{b}{a})\) if you are in quadrant I or IV because that is the range of arctangent, but you are in quadrant 2

OpenStudy (anonymous):



OpenStudy (anonymous):

\(\theta\) is not unique, you can go around the circle again and again to come to a coterminal angle, but the one you found, \(-.519146\) is in the wrong place, it is in quadrant 4

OpenStudy (anonymous):

easy enough to fix though, add \(\pi\) to your answer

OpenStudy (anonymous):

is this more or less clear? arctangent only returns answers in the interval \([-\frac{\pi}{2},\frac{\pi}{2}]\) and you are not in that interval